OpenStudy (anonymous):

How to find total area of the region between curve and x-axis??? y=2x+7; 1≤x≤5 and y=(4x)/(1+x^2), interval -1≤x≤1

OpenStudy (mathmate):

For f(x)=2x+7, and \(as\ long\ as\ f(x)\ge0\ in\ the\ given\ interval\ [1,5]\), the area is given by \(\int_1^5 f(x)dx\).

OpenStudy (anonymous):

what about the secondn one?

OpenStudy (mathmate):

@kevst3r Can you tell me if f(x)=(4x)/(1+x^2) is even or odd?

OpenStudy (anonymous):

odd

OpenStudy (mathmate):

For an odd function, f(-x)=-f(x). IF f(x) in the interval [0,1] is >= 0, then f(x) in the interval [-1,0] is <=0. Thus the total area between the curve and the x-axis is equal to \(\int_0^1 f(x)dx - \int_{-1}^0 f(x)dx\), or simply using the properties of odd functions: \(\int_0^1 f(x)dx - \int_{-1}^0 f(x)dx = 2\int_0^1 f(x)dx\) 

OpenStudy (anonymous):

the answer is \[(1/3)\ln(lnx)+C\]

OpenStudy (anonymous):

@mathmate so how do ig et that

OpenStudy (mathmate):

Integrate f(x)=4x/(1+x^2) Note that (except for constants), the numerator is the derivative of the denominator. Use then substitution s=(1+x^2), ds=2x dx, or 2ds = 4x dx Make the appropriate substitutions, \(\int f(x)dx = \int \frac{4xdx}{1+x62}=\int 2ds/s = 2log(s)=2 log(1+x^2)+C\) but C disappears (cancels out) when you evaluate definite integrals: \(2\int_0^1 f(x)dx = 2\int_0^1 \frac{4xdx}{1+x^2}dx=2[log(1+x^2)]_0^1=2*log(2)=4log(2)\) where log(x) is natural log (usually written as ln(x) in highschool).

OpenStudy (mathmate):

Typo: \(2\int_0^1 f(x)dx = 2\int_0^1 \frac{4xdx}{1+x^2}dx=2[2log(1+x^2)]_0^1=2*2log(2)=4log(2)\)